OpenStudy (anonymous):

Evaluate the polynomial 6x-7 for 1/2n^2 - 1/2n and y=2 A. 15 B. -20 C. 4 D. -16

OpenStudy (vishal_kothari):

what is y and n here?

OpenStudy (anonymous):

thats the way its written, i could'nt understand it myself

OpenStudy (anonymous):

\[f(x)=6x-7\rightarrow f(1/2n^2-1/2n)=3/n^2-3/n-7\]

OpenStudy (anonymous):

\[f(2)=6(2)-7=5\]

OpenStudy (anonymous):

I think it is just a straightforward substitution.

OpenStudy (anonymous):

?